OpenStudy (anonymous):

Multiply and then simplify the radical. 5 square root of 10 times 3 square root of 8 A. 15 square root of 80 B. 15 square root of 18 C. 45 square root of 5 D. 60 square root of 5

OpenStudy (fretje):

5* sqrt(10) * 3 * sqrt(8) = 3 * 5 * sqrt (10 * 8) = 15 * sqrt (80)

OpenStudy (fretje):

so answer A